# KEPLA: A Knowledge-Enhanced Deep Learning Framework for Accurate Protein-Ligand Binding Affinity Prediction

### v1

@@ Abstract
Accurate prediction of protein-ligand binding affinity is critical for drug discovery. While recent deep learning approaches have demonstrated promising results, they often rely solely on structural features, overlooking valuable biochemical knowledge associated with binding affinity. To address this limitation, we propose KEPLA, a novel deep learning framework that explicitly integrates prior knowledge from Gene Ontology and ligand properties of proteins and ligands to enhance prediction performance. KEPLA takes protein sequences and ligand molecular graphs as input and optimizes two complementary objectives: (1) aligning global representations with knowledge graph relations to capture domain-specific biochemical insights, and (2) leveraging cross-attention between local representations to construct fine-grained joint embeddings for prediction. Experiments on two benchmark datasets across both in-domain and cross-domain scenarios demonstrate that KEPLA consistently outperforms state-of-the-art baselines. Furthermore, interpretability analyses based on knowledge graph relations and cross-attention maps provide valuable insights into the underlying predictive mechanisms.
@@ Introduction
The prediction of protein-ligand binding affinity (PLA) has been widely considered as one of the most important tasks in computational drug discovery [1]. Here, ligands are usually drug candidates, including small molecules and biologics, which can interact with proteins as agonists or inhibitors in biological processes to cure diseases. Given a protein, we want to understand how well a drug molecule (called a ligand) can interact with this protein. The strength of interaction between them can be quantified as a numerical score (called the binding affinity), which potentially determines whether a ligand can effectively influence the protein (for example, to inactivate a protein to cure a disease). Although experimental methods [2] can measure binding affinity, these biological tests are laborious and time-consuming. Thus, in silico predictive approaches have become increasingly necessary and received widespread attention [3]. With in silico approaches, binding affinities can be quickly predicted in the early stage of drug discovery. The predicted binding affinities can help rank drug candidates and prioritize the appropriate ones for more precise subsequent tests to accelerate the drug screening process, instead of applying costly biological methods directly to screen numerous candidate molecules [4].
@@
Recently, along with their success in other areas, deep learning approaches have rapidly progressed for computational PLA prediction. The scientific community has primarily divided these deep learning approaches into interaction-based and interaction-free methodologies, which are distinguished by whether they rely on physical interactions to inform predictions. Interaction-based approaches make predictions based on the known three-dimensional (3D) structures of complexes and the physical interactions between proteins and ligands [5, 6]. As the number of biological targets with available 3D structures is limited, interaction-based approaches are less useful than interaction-free ones when faced with unknown protein-ligand complexes. In contrast, without focusing on direct physical interactions, interaction-free approaches infer binding affinity from protein and ligand lower-dimensional data, such as amino acid sequence and simplified molecular input line entry system (SMILES) [7]. Regarding PLA prediction as a regression task, interaction-free approaches first feed the protein and ligand inputs into different encoders such as convolutional neural network (CNN) [8, 9], graph neural network (GNN) [10, 11], and Transformer [12, 13, 14]. With these advanced deep learning techniques, data-driven representations can be automatically learned from large-scale protein and ligand data. Subsequently, the binding affinities between proteins and ligands are predicted through a decoder that incorporates and processes their representations. Obviously, interaction-free methods have wider applications, as they take 1D or 2D structural information from proteins and ligands as input and thus have fewer restrictions on data.
@@
Despite their promising results, existing interaction-free prediction models still suffer from the following limitations: relying solely on structural information creates performance bottlenecks, and the predictions lack scientific interpretability. To address these issues, we propose that computational models should incorporate biochemical factual knowledge when encoding proteins and ligands. Such knowledge contains rich external information beneficial for PLA prediction, which can enhance predictive performance while serving as scientific evidence to provide interpretability for the prediction results. For example, proteins’ Gene Ontology (GO) [15] encompasses knowledge about their molecular functions, biological processes, and cellular components. This information describes the biological characteristics of proteins and is closely related to protein-ligand binding. For instance, if a protein’s molecular function is identified as ATP binding, this may suggest that the protein could exhibit a high affinity for ATP-like ligands. Biological processes such as signal transduction might indicate the protein’s involvement in specific pathways and potential interactions with other molecules. Cellular components such as the cell membrane could influence the accessibility of the ligand. Furthermore, ligand properties (LP) reveal knowledge from molecular descriptors and chemical features. For example, hydrogen bond donors (e.g., amino groups, phenolic hydroxyl groups) and hydrogen bond acceptors (e.g., oxygen, nitrogen) in ligand molecules can form hydrogen bonds with proteins, playing crucial roles in protein-ligand binding, and thus understanding their specific number is essential for predicting PLA.
@@
However, integrating such domain knowledge introduces two key challenges: (1) how to organize the diverse and complex knowledge into a form that the model can easily process; (2) how to enable the model to sufficiently capture the organized knowledge. To address the first challenge, knowledge graphs (KGs) [16] are expected to be the optimal solution due to their proven capabilities in knowledge tasks. KGs from GO and LP can explicitly describe extensive biochemical facts via the form of entity-relation-entity triples. Correspondingly, KG embedding approaches [17] can efficiently embed proteins, ligands, and other knowledge entities into continuous vectors of entities and relations. To address the second challenge, the seemingly simplest approach is to use the embeddings of proteins and ligands from the KG, together with the representations generated by the structural information encoder for PLA prediction. However, this approach is ineffective for knowledge-deficient proteins or ligands, since KG embedding cannot work like structural information encoders when only their amino acid sequence or SMILES string is available. Therefore, incorporating the KG to enhance a structural information encoder is an intuitively better strategy that can offer the best of both worlds. We build a bridge between structural information encoding and KG embedding, using a multiple objective learning approach to ensure that knowledge is injected during the encoding process from structural information to representations.
@@
In this paper, we propose a knowledge-enhanced protein-ligand binding affinity prediction (KEPLA) model, which is the first general framework to integrate biochemical factual knowledge into protein-ligand binding affinity prediction. Specifically, the main framework of the model still employs an encoder-decoder architecture. It employs hybrid encoders separately based on evolutionary sequence model (ESM) [18] agent and graph convolutional networks (GCNs) [19] to encode 1D amino acid sequences and 2D molecular graphs into protein and ligand local representations. To ensure the encoders capture the relevant biochemical knowledge and exploit it for PLA prediction, we jointly utilize the encoded protein and ligand representations for KG embedding and PLA prediction regression objectives during model training. For the KG embedding objective, the encoded protein and ligand global representations are directly used as the entity embeddings of our constructed KG, which covers protein-GO and ligand-LP knowledge. Then they are optimized along with learnable embedding parameters of other entities and relations in the KG following vanilla KG embedding approaches. For the PLA objective, the encoded local representations are fed into a pairwise interaction module that consists of a cross attention network [20] to learn protein-ligand joint representations. The joint representations are then decoded by a multilayer perceptron (MLP) decoder to make a PLA prediction. By jointly minimizing the KG embedding and the PLA prediction losses, our model takes into account the relevant biochemical knowledge while making PLA predictions and thus can be enhanced. We conduct a comprehensive experimental comparison against state-of-the-art PLA prediction methods in in-domain and cross-domain settings. Experimental results on both datasets demonstrate that our method improves prediction accuracy by 5.28% and 12.42% compared to state-of-the-art methods (using RMSE as the metric), respectively, while also providing more comprehensive interpretable insights for the predictions.
@@
To summarize, our contributions are as follows: (1) We propose KEPLA, the first knowledge-enhanced PLA predicting model that incorporates a KG with the interaction-free framework and brings promising performance improvements. (2) By learning with the KG and attention mechanism, KEPLA gives an interpretable prediction with cross attention weights and relevant knowledge instead of black-box results. (3) We construct and release a novel KG dataset based on PDBbind, promoting the research on PLA prediction.
@@ Results
The proposed KEPLA framework is illustrated in Figure 1. Given an input protein-ligand pair, we first encode the protein amino acid sequence and the ligand molecular graph separately using ESM and GCN blocks, respectively. To incorporate relevant biochemical knowledge into the representation learning process, the encoded representations are further embedded into a KG and optimized with a KG embedding objective. Simultaneously, a cross attention module is employed to capture fine-grained interactions between the protein and ligand representations. This module includes a local interaction mapping step followed by a cross attention computation step, which together generate a joint protein-ligand representation. Finally, a fully connected MLP decoder produces a predictive score representing the binding affinity. Details of each module in the KEPLA model, along with the training procedure, are provided in the Methods section.
Figure 1: Overview of the KEPLA framework. (a) Strategy of existing PLA prediction models. (b) Strategy of our knowledge-enhanced model. The key difference lies in the incorporation of a biochemical KG, which enhances both protein and ligand encoders. (c) Detailed architecture of KEPLA. The model employs two separate encoders: an ESM-based encoder processes the amino acid sequence, and a GCN-based encoder handles the SMILES string. These generate both local and global representations for proteins and ligands. The global representations are used in the KG embedding module, while the local representations are input into a cross attention module for PLA prediction. (d) The cross attention module. Here, an interaction map is computed via the multiplication of protein and ligand local representations, which serves as a similarity matrix to derive cross attention. The cross attention enables fine-grained fusion of protein and ligand local representations, resulting in a joint representation. This joint representation is passed through an MLP decoder to predict binding affinity. (e) The KG embedding module. Here, linear transformations project the global protein and ligand representations into a semantic space. Combined with entity and relation embeddings from the KG, a score function evaluates the plausibility of triplet relationships, which forms the objective of this module.
@@
We evaluate prediction performance on two public datasets, PDBbind [21] and CSAR-HiQ [22], using two distinct data split strategies to simulate in-domain and cross-domain scenarios. For in-domain evaluation, following [23], we use the refined set of PDBbind as the primary training data, as there is substantial overlap between the full general set and CSAR-HiQ dataset. We randomly split the protein-ligand complexes in the refined set into training and validation sets at a 9:1 ratio. For testing, we use the core set and the CSAR-HiQ dataset after removing any overlapping complexes from the refined set. For cross-domain evaluation, we propose a clustering-based pair split strategy to simulate domain shift. We conduct cross-domain evaluation on the larger-scale PDBbind datasets. Specifically, we first use the single-linkage algorithm to cluster ligands and proteins by ECFP4 (extended connectivity fingerprint, up to four bonds) [24] and PSC (pseudo-amino acid composition) [25], respectively. After that, we randomly select 60% ligand clusters and 60% protein clusters from the clustering result, and consider all protein-ligand pairs between the selected ligands and proteins as source domain data. All the pairs between ligands and proteins in the remaining clusters are considered to be target domain data. Detailed clustering implementation is provided in the Supplementary Information, section 1. Under this split, source and target domains are disjoint and follow different distributions. In line with standard domain adaptation settings, we use all labeled source domain data and 80% of the unlabeled target domain data for training, where this 80% also serves as the validation set. The remaining 20% of labeled target data is used for testing. Compared to in-domain random splits, this cross-domain setup poses a greater challenge but better reflects real-world generalization scenarios in drug discovery.
@@
Table 1 presents the quantitative results. We report the mean and standard deviation of four evaluation metrics across five independent runs. When comparing model categories, we observe that interaction-based models generally outperform interaction-free models, highlighting the advantage of incorporating structural information and physical interactions in PLA prediction. Despite being an interaction-free method, our proposed KEPLA achieves the best performance across all experimental settings on both datasets, surpassing even all interaction-based baselines. Specifically, KEPLA improves RMSE by 5.28% on the PDBbind dataset and by 12.42% on the CSAR-HiQ dataset compared to the best-performing baseline GIANT. These results strongly support KEPLA’s superior predictive capability. Considering KEPLA requires only the molecular graph of ligand compounds and primary sequences of target proteins as input like most interaction-free methods, the outstanding performance benefits from knowledge-enhanced representation learning and local interaction modeling.
@@
00footnotetext: Results are reported as mean (standard deviation). For each dataset and metric, the best performance is highlighted in bold, and the second-best is underlined.
@@
In-domain classification under a random split setting is relatively easier and holds limited practical significance. Therefore, we further investigate the more realistic and challenging task of cross-domain PLA prediction, where the training and test data follow different distributions. To simulate this scenario, we partition the original PDBbind dataset into source and target domains using the clustering-based pair split strategy. As a complementary evaluation, we also perform experiments using a cold pair split on the PDBbind dataset. Specifically, we randomly select 70% of the proteins and ligands and collect all associated protein-ligand pairs as the training set. The remaining 30% of proteins and ligands are used to form the validation and test sets in a 3:7 ratio. This cold pair split strategy ensures that none of the proteins or ligands in the test set have been seen during training, providing a stringent test of model generalization.
The performance evaluation on the PDBbind dataset under both clustering-based and cold pair split strategies is shown in Figure 2. It is important to note that these experiments are restricted to interaction-free methods, as structural interaction information is unavailable in the simulated cross-domain and cold scenarios. Compared to the earlier in-domain results, the performance of all interaction-free PLA models drops significantly as shown in Figure 2a. This decline is primarily attributed to the reduced information overlap between the training and test sets, as well as the absence of sufficiently similar proteins and ligands to serve as references. Moreover, Figure 2b highlights a noticeable performance drop when moving from random split to cold pair split, emphasizing that prediction accuracy on unseen test data cannot solely rely on previously encountered ligand or protein features.
@@
Figure 3: Ablation study in terms of RMSE and R on the PDBbind and CSAR-HiQ datasets under random split (statistics over five independent runs). (a) Performance of different variants of the KG module. (b) Performance of different variants of the cross-attention module. The vertical bars represent the mean values, while the black lines denote standard deviations (error bars). Dots indicate the performance scores from each individual run.
@@
We conduct an ablation study to investigate the contributions of two key components in KEPLA: the KG and the cross attention modules. The results of various KEPLA variants are presented in Figure 3. To evaluate the impact of KG enhancement, we consider three variants of KEPLA with different configurations: KEPLA using only the protein-side KG, KEPLA using only the ligand-side KG, and KEPLA with no KG enhancement. As shown in Figure 3a, the model’s strong performance is clearly supported by the incorporation of external knowledge, validating the effectiveness of KG enhancement. Notably, the protein-side KG contributes more significantly to affinity prediction than the ligand-side counterpart. A likely explanation is that the protein KG contains a greater number of entities and richer relational information, offering more comprehensive context for representation learning.
@@
To further assess the effectiveness of the cross attention mechanism, we evaluate three KEPLA variants that differ in how they compute joint representations between proteins and ligands: one-side ligand attention, one-side protein attention, and linear concatenation. In these variants, the cross attention module in KEPLA is replaced with either ligand-side or protein-side attention. The one-side attention variants adopt the neural attention mechanism proposed in [40], where the joint representation is derived from the interaction between a ligand vector and a protein subsequence matrix. The linear concatenation variant, on the other hand, simply concatenates the ligand and protein vector representations following a max-pooling operation. As shown in Figure 3b, the results confirm that the cross attention mechanism is both effective and essential, as it enables the model to capture fine-grained interactions crucial for accurate PLA prediction. Among the one-side attention variants, ligand-side attention yields better performance. This may be attributed to the fact that ligand-side attention operates on local representations derived from the ligand’s 2D molecular structure, which tend to be more spatially precise compared to the 1D sequence-based representations of proteins.
@@
Figure 4: Visualization of ligands and binding pockets for structural-level interpretability study. (a) Interpretability of co-crystallized ligands. The upper section of each panel displays the 2D structures of ligands, with atoms highlighted in orange to indicate those predicted to contribute to protein binding. All 2D structures were visualized using RDKit. The lower section of each panel presents protein-ligand interaction maps derived from the corresponding crystal structures. (b) Interpretability of binding pocket structures. The 3D representations show protein-ligand binding pockets, with correctly predicted binding-site residues highlighted in orange and the corresponding ligands shown in cyan. Remaining amino acid residues, secondary structure elements, and surface maps are shown in grey. All protein-ligand interaction maps and 3D visualizations of X-ray crystal structures were generated using the Molecular Operating Environment (MOE) software.

### v2

@@ Abstract
Accurate prediction of protein-ligand binding affinity is critical for drug discovery. While recent deep learning approaches have demonstrated promising results, they often rely solely on structural features of proteins and ligands, overlooking their valuable biochemical knowledge associated with binding affinity. To address this limitation, we propose KEPLA, a novel deep learning framework that explicitly integrates prior knowledge from Gene Ontology and ligand properties to enhance prediction performance. KEPLA takes protein sequences and ligand molecular graphs as input and optimizes two complementary objectives: (1) aligning global representations with knowledge graph relations to capture domain-specific biochemical insights, and (2) leveraging cross attention between local representations to construct fine-grained joint embeddings for prediction. Experiments on two benchmark datasets across both in-domain and cross-domain scenarios demonstrate that KEPLA consistently outperforms state-of-the-art baselines. Furthermore, interpretability analyses based on knowledge graph relations and cross attention maps provide valuable insights into the underlying predictive mechanisms.
@@ Introduction
The prediction of protein-ligand binding affinity (PLA) has been widely considered as one of the most important tasks in computational drug discovery [1]. Here, ligands are usually drug candidates, including small molecules and biologics, which can interact with proteins as agonists or inhibitors in biological processes to cure diseases. Given a protein, we want to understand how well a ligand can interact with this protein. The strength of interaction between them can be quantified as a numerical score (called the binding affinity), which potentially determines whether a ligand can effectively influence the protein (for example, to inactivate a protein to cure a disease). Although experimental methods [2] can measure binding affinity, these biological tests are laborious and time-consuming. Thus, in silico predictive approaches have become increasingly necessary and received widespread attention [3]. With in silico approaches, binding affinities can be quickly predicted in the early stage of drug discovery. The predicted binding affinities can help rank drug candidates and prioritize the appropriate ones for more precise subsequent tests to accelerate the drug screening process, instead of applying costly biological methods directly to screen the entire pool of numerous candidate molecules [4].
@@
Recently, along with their success in other areas, deep learning approaches have rapidly progressed for computational PLA prediction [5]. The scientific community has primarily divided these deep learning approaches into interaction-based and interaction-free methodologies, which are distinguished by whether they rely on physical interactions to inform predictions. Interaction-based approaches make predictions based on the known three-dimensional (3D) structures of complexes and the physical interactions between proteins and ligands [6, 7, 8]. As the number of biological targets with available 3D structures is limited, interaction-based approaches are less useful than interaction-free ones when faced with unknown protein-ligand complexes. In contrast, without relying on direct physical interactions, interaction-free approaches infer binding affinity from protein and ligand lower-dimensional data, such as amino acid sequence and simplified molecular input line entry system (SMILES) string [9]. Regarding PLA prediction as a regression task, interaction-free approaches first feed the protein and ligand inputs into different encoders such as convolutional neural network (CNN) [10, 11], graph neural network (GNN) [12, 13], and Transformer [14, 15, 16]. With these advanced deep learning techniques, data-driven representations can be automatically learned from large-scale protein and ligand data. Subsequently, the binding affinities between proteins and ligands are predicted through a decoder that incorporates and processes their representations. Obviously, interaction-free methods have wider applications, as they take 1D or 2D structural information from proteins and ligands as input and thus have fewer restrictions on data.
@@
Despite their promising results, existing interaction-free models still suffer from two key limitations: (1) their reliance solely on structural information creates performance bottlenecks, and (2) their predictions often lack scientific interpretability. To overcome these limitations, we argue that interaction-free models should incorporate biochemical factual knowledge during the encoding of proteins and ligands. Such knowledge provides rich external information that not only enhances predictive accuracy but also offers scientific grounding for interpretability. For proteins, Gene Ontology (GO) [17] annotations describe their molecular functions, biological processes, and cellular components, which are highly relevant to protein-ligand binding. For instance, if a protein’s molecular function includes ATP binding, it may exhibit high affinity for ATP-like ligands. Involvement in biological processes such as signal transduction can suggest potential molecular interactions or pathway associations. Similarly, cellular component annotations like cell membrane can inform ligand accessibility. For ligands, knowledge can be derived from molecular descriptors and chemical features. Ligand properties (LP) such as the number of hydrogen bond donors (e.g., amino groups, phenolic hydroxyl groups) and acceptors (e.g., oxygen, nitrogen atoms) are crucial for forming hydrogen bonds with proteins. These interactions play a pivotal role in protein-ligand binding, making such properties essential for accurate PLA prediction.
@@
However, integrating such domain knowledge presents two key challenges: (1) how to organize the diverse and complex knowledge into a form that the model can easily process, and (2) how to enable the model to sufficiently capture the organized knowledge. To address the first challenge, knowledge graphs (KGs) [18] are expected to be the optimal solution due to their proven capabilities in knowledge tasks. KGs constructed from GO and LP can systematically describe biochemical knowledge of proteins and ligands via the form of entity-relation-entity triples. Correspondingly, KG embedding techniques [19] can map proteins, ligands, and related entities into continuous vector spaces, preserving both semantic and relational information. For the second challenge, a straightforward approach is to directly use the KG-derived embeddings of proteins and ligands, alongside structural representations, for PLA prediction. However, this approach falls short for knowledge-sparse entities, as KG embeddings cannot substitute structural encoders when only amino acid sequences or SMILES strings are available. Therefore, incorporating the KG to enhance the structural information encoder is a more effective strategy that can offer the best of both worlds. To this end, we propose a joint learning framework that bridges structural encoding and KG embedding through multi-objective learning, ensuring that biochemical knowledge is effectively injected during the representation learning process.
@@
In this paper, we propose a knowledge-enhanced protein-ligand binding affinity prediction (KEPLA) model, which is the first general framework to achieve deep integration of biochemical factual knowledge into the task of PLA prediction. The overall architecture of KEPLA follows an encoder-decoder paradigm. Specifically, it employs hybrid encoders—evolutionary sequence model (ESM) [20] for 1D amino acid sequences and graph convolutional networks (GCNs) [21] for 2D molecular graphs— to generate both global and local representations of proteins and ligands. To ensure the encoders effectively capture and leverage biochemical knowledge for the task, KEPLA jointly trains the encoded protein and ligand representations on two objectives: KG embedding and PLA prediction. For the KG embedding objective, the encoded global representations are directly used as the entity embeddings of our constructed KG, which covers protein-GO and ligand-LP knowledge. Then they are optimized alongside other learnable entity and relation embeddings using standard KG embedding techniques. For the PLA prediction objective, the encoded local representations are processed by a pairwise interaction module, which uses a cross attention network [22] to derive protein-ligand joint representations. The joint representations are then decoded by a multilayer perceptron (MLP) decoder to produce the final affinity prediction. By jointly minimizing the losses of both objectives, our model takes into account the relevant biochemical knowledge while making PLA predictions and thus can be enhanced. We conduct extensive experiments under both in-domain and cross-domain settings, comparing KEPLA with state-of-the-art methods. Results demonstrate that KEPLA achieves significant performance gains, improving RMSE metric by 5.28% and 12.42% on two benchmark datasets, respectively. Moreover, KEPLA provides more interpretable insights into the prediction through its knowledge-grounded framework.
@@
To summarize, our contributions are as follows: (1) We propose KEPLA, the first knowledge-enhanced PLA predicting model that incorporates a KG with the interaction-free framework, achieving notable performance improvements. (2) By leveraging the KG and cross attention mechanism, KEPLA provides an interpretable prediction through relevant knowledge and attention visualization, moving beyond traditional black-box outputs. (3) We construct and release a novel KG dataset based on PDBbind, promoting the research on PLA prediction.
@@ Results
The proposed KEPLA framework is illustrated in Figure 1. Given an input protein-ligand pair, we first encode the protein amino acid sequence and the ligand molecular graph separately using ESM and GCN blocks, respectively. To incorporate relevant biochemical knowledge into the representation learning process, the encoded representations are further embedded into a KG module and optimized with the KG embedding objective. Simultaneously, a cross attention module is employed to capture fine-grained interactions between the protein and ligand representations. This module includes a local interaction mapping step followed by a cross attention computation step, which together generate a joint protein-ligand representation. Finally, a fully connected MLP decoder produces a predictive score representing the binding affinity. Details of each module in the KEPLA model, along with the training procedure, are provided in the Methods section.
Figure 1: Overview of the KEPLA framework. (a) Strategy of existing PLA prediction models. (b) Strategy of our KEPLA model. The key difference lies in the incorporation of a biochemical KG, which enhances both protein and ligand encoding. (c) Detailed architecture of KEPLA. The model employs two separate encoders: an ESM-based encoder processes the amino acid sequence, and a GCN-based encoder handles the SMILES string. These generate both local and global representations for proteins and ligands. The global representations are used in the KG embedding module, while the local representations are input into a cross attention module for PLA prediction. (d) The cross attention module. Here, an interaction map is computed via the multiplication of protein and ligand local representations, which serves as a similarity matrix to derive cross attention. The cross attention enables fine-grained fusion of protein and ligand local representations, resulting in a joint representation. This joint representation is passed through an MLP decoder to predict binding affinity. (e) The KG embedding module. Here, linear transformations project the global protein and ligand representations into a semantic space. Combined with entity and relation embeddings from the KG, a score function evaluates the plausibility of triplets, which forms the objective of this module.
@@
We evaluate prediction performance on two public datasets, PDBbind [23] and CSAR-HiQ [24], using two distinct data split strategies to simulate in-domain and cross-domain scenarios. For in-domain evaluation, following [25], we use the refined set of PDBbind as the primary training data, as there is substantial overlap between the full general set and CSAR-HiQ dataset. We randomly split the protein-ligand complexes in the refined set into training and validation sets at a 9:1 ratio. For testing, we use the PDBbind core set and the CSAR-HiQ dataset after removing any overlapping complexes from the refined set. For cross-domain evaluation, we propose a clustering-based pair split strategy to simulate domain shift. We conduct cross-domain evaluation on the larger-scale PDBbind datasets. Specifically, we first use the single-linkage algorithm to cluster proteins and ligands by PSC (pseudo-amino acid composition) [26] and ECFP4 (extended connectivity fingerprint, up to four bonds) [27], respectively. After that, we randomly select 60% protein clusters and 60% ligand clusters from the clustering result, and consider all protein-ligand pairs between the selected ligands and proteins as source domain data. All the pairs between proteins and ligands in the remaining clusters are considered to be target domain data. Detailed clustering implementation is provided in the Supplementary Information. Under this split, source and target domains are disjoint and follow different distributions. In line with standard domain adaptation settings, we use all labeled source domain data and 80% of the unlabeled target domain data for training, where this 80% also serves as the validation set. The remaining 20% of labeled target data is used for testing. Compared to in-domain random splits, this cross-domain setup poses a greater challenge but better reflects real-world generalization scenarios in drug discovery.
@@
Table 1 presents the quantitative results. We report the mean and standard deviation of four evaluation metrics across five independent runs. When comparing model categories, we observe that interaction-based models generally outperform interaction-free models, highlighting the advantage of incorporating 3D structural information and physical interactions in PLA prediction. Despite being an interaction-free method, our proposed KEPLA achieves the best performance across all experimental settings on both datasets, surpassing even all interaction-based baselines. Specifically, KEPLA improves RMSE by 5.28% on the PDBbind dataset and by 12.42% on the CSAR-HiQ dataset compared to the best-performing baseline GIANT. These results strongly support KEPLA’s superior predictive capability. Considering KEPLA requires only the molecular graph of ligand compounds and primary sequences of target proteins as input like most interaction-free methods, the outstanding performance benefits from knowledge-enhanced representation learning and local interaction modeling.
@@
  • Results are reported as mean (standard deviation). For each dataset and metric, the best performance is highlighted in bold, and the second-best is underlined.
@@
In-domain prediction under a random split setting is relatively easier and holds limited practical significance. Therefore, we further investigate the more realistic and challenging task of cross-domain PLA prediction, where the training and test data follow different distributions. To simulate this scenario, we partition the original PDBbind dataset into source and target domains using the clustering-based pair split strategy. As a complementary evaluation, we also perform experiments using a cold pair split on the PDBbind dataset. Specifically, we randomly select 70% of the proteins and ligands and collect all associated protein-ligand pairs as the training set. The remaining 30% of proteins and ligands are used to form the validation and test sets in a 3:7 ratio. This cold pair split strategy ensures that none of the proteins or ligands in the test set have been seen during training, providing a stringent test of model generalization.
The performance evaluation on the PDBbind dataset under both clustering-based and cold pair split strategies is shown in Figure 2. It is important to note that these experiments are restricted to interaction-free methods, as structural interaction information is unavailable in the simulated cross-domain and cold scenarios. Compared to the earlier in-domain results, the performance of all interaction-free PLA models drops significantly as shown in Figure 2a. This decline is primarily attributed to the reduced information overlap between the training and test sets, as well as the absence of sufficiently similar proteins and ligands to serve as references. Moreover, Figure 2b highlights a noticeable performance drop when moving from random split to cold pair split, emphasizing that prediction accuracy on unseen test data cannot solely rely on previously encountered protein or ligand features.
@@
Figure 3: Ablation study in terms of RMSE and R on the PDBbind and CSAR-HiQ datasets under random split (statistics over five independent runs). (a) Performance of different variants of the KG module. (b) Performance of different variants of the cross attention module. The vertical bars represent the mean values, while the black lines denote standard deviations (error bars). Dots indicate the performance scores from each individual run.
@@
We conduct an ablation study to investigate the contributions of two key components in KEPLA: the KG and the cross attention modules. The results of various KEPLA variants are presented in Figure 3. To evaluate the impact of KG enhancement, we consider three variants of KEPLA with different configurations: KEPLA using only the protein-side KG, KEPLA using only the ligand-side KG, and KEPLA without KG enhancement. As shown in Figure 3a, the model’s strong performance is clearly supported by the incorporation of external knowledge, validating the effectiveness of KG enhancement. Notably, the protein-side KG contributes more significantly to affinity prediction than the ligand-side counterpart. A likely explanation is that the protein KG contains a greater number of entities and richer relational information, offering more comprehensive context for representation learning.
@@
To further assess the effectiveness of the cross attention mechanism, we evaluate three KEPLA variants that differ in how they compute joint representations between proteins and ligands: one-side protein attention, one-side ligand attention, and linear concatenation. In these variants, the cross attention module in KEPLA is replaced with either ligand-side or protein-side attention. The one-side attention variants adopt the neural attention mechanism proposed in [42], where the joint representation is derived from the interaction between a ligand vector and a protein fragment matrix. The linear concatenation variant, on the other hand, simply concatenates the protein and ligand vector representations following a max-pooling operation. As shown in Figure 3b, the results confirm that the cross attention mechanism is both effective and essential, as it enables the model to capture fine-grained interactions crucial for accurate PLA prediction. Among the one-side attention variants, ligand-side attention yields better performance. This may be attributed to the fact that ligand-side attention operates on local representations derived from the ligand’s 2D molecular structure, which tend to be more spatially precise compared to the 1D sequence-based representations of proteins.
@@
Figure 4: Visualization of ligands and binding pockets for structural-level interpretability study. (a) Interpretability of co-crystallized ligands. The upper section of each panel displays the 2D structures of ligands, with atoms highlighted in orange to indicate those predicted to contribute to protein binding. All 2D structures are visualized using RDKit. The lower section of each panel presents protein-ligand interaction maps derived from the corresponding crystal structures. (b) Interpretability of binding pocket structures. The 3D representations show protein-ligand binding pockets, with correctly predicted binding-site residues highlighted in orange and the corresponding ligands shown in cyan. Remaining amino acid residues, secondary structure elements, and surface maps are shown in grey. All protein-ligand interaction maps and 3D visualizations of X-ray crystal structures are generated using the Molecular Operating Environment (MOE) software.